Schedule of Cash Payments

EastGate Physical Therapy Inc. is planning its cash payments for operations for the first quarter (January–March). The Accrued Expenses Payable balance on January 1 is $25,800. The budgeted expenses for the next three months are as follows:

January February March
Salaries $59,300 $72,200 $80,000
Utilities 4,900 5,400 6,500
Other operating expenses 45,700 49,800 54,800
Total $109,900 $127,400 $141,300

Other operating expenses include $3,200 of monthly depreciation expense and $700 of monthly insurance expense that was prepaid for the year on May 1 of the previous year. Of the remaining expenses, 65% are paid in the month in which they are incurred, with the remainder paid in the following month. The Accrued Expenses Payable balance on January 1 relates to the expenses incurred in December.

Prepare a schedule of cash payments for operations for January, February, and March.

EastGate Physical Therapy Inc.
Schedule of Cash Payments for Operations
For the Three Months Ending March 30
January February March
Payments of prior month's expense $ $ $
Payments of current month's expense      
Total payments $ $ $   
Answer #1

Prepare a schedule of cash payments for operations for January, February, and March.

(1)

Total expense = $109900

Depreciation = $3200

Insurance = $700

Paid percentage = 35%

Paid for prior month exp = ($109900 - $3200 - $700) * 35% = $37100

(2)

Total expense = $127400

Depreciation = $3200

Insurance = $700

Paid percentage = 35%

Paid for prior month exp = ($127400 - $3200 - $700) * 35% = $43225

(3)

37100/35*65 = $68900

(4) $43225/35*65 = $80275

(5)

Total expense = $141300

Depreciation = $3200

Insurance = $700

Paid percentage = 35%

Paid for prior month exp = ($141300 - $3200 - $700) * 65% = $89310

EastGate Physical Therapy Inc.
Schedule of Cash Payments for Operations
For the Three Months Ending March 30
January February March
Payments of prior month's expense $25800 $37100 (1) $43225 (2)
Payments of current month's expense $68900 (3) $80275 (4)

$89310 (5)

Total payments $94700 $117375 $132535